OpenStudy (anonymous):

Determine the value or values of d that complete the square for the expression. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.) x2 + dx + 36 x2 + dx + (49/4) x2 + 7x + d

OpenStudy (anonymous):

plz help me!!!!

OpenStudy (anonymous):

12,7,7/2.... the first is just (x+6)^2 and the 2nd is (x+7/2)^2 and the third is just (x+7/2)^2 also

OpenStudy (anonymous):

12,7,(49/4) sorry

OpenStudy (anonymous):

value of \(d\) you want first one, \(36=6^2\) and \(2\times 6=12\) so that is your \(d\)

OpenStudy (anonymous):

thanks